OpenStudy (kyanda17):

If the value of 2x3 is 2, then what is the value of x? If the value of 2x3 is 2, then what is the value of x? a.1/3 b.1/2 c.1

OpenStudy (hybrik):

Set an equation: 2(x)3= 2

OpenStudy (hybrik):

Or (2x)3=2

OpenStudy (hybrik):

Divide both sides by 3: 2x=2/3

OpenStudy (hybrik):

When divide both sides by 2, x=2/6 , then simplify, x=1/3

OpenStudy (kyanda17):

thank you

OpenStudy (michele_laino):

I think that the equation is like below: \[\Large 2{x^3} = 2\]

OpenStudy (michele_laino):

am I right?

OpenStudy (kyanda17):

yes

OpenStudy (hybrik):

oh, make sure to put it as 2x^3...